OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
BHS < > MTL
Scheduled Maintenance Report for Network & Infrastructure
Completed
We have got an incident on 4 links between the datacenter of Beauharnois and the PoP of Montreal.

We raised a ticket to our provider

Update(s):

Date: 2018-04-13 00:19:43 UTC
Back to normal

Date: 2018-04-12 01:05:02 UTC
The reparation has been done.

We are waiting for the provider to give us the status and the green light regarding this fiber cut

Date: 2018-04-12 01:03:20 UTC
There is a fiber cut at 22 kms of the Beauharnois datacenter.
The team of the provider is in place to repair the issue. No ETA for the moment.


Date: 2018-04-11 01:18:29 UTC
We are going to shift ROGERS to Chicago in order to don't overload the links in Toronto

Date: 2018-04-11 00:27:11 UTC
We've shifted Level3 too

Date: 2018-04-11 00:11:36 UTC
We have shifted BELL to Toronto

http://weathermap.ovh.net/#usa
Posted Apr 11, 2018 - 00:10 UTC